The first ionization energies of cesium and gold are 3.89 and 9.23 eV, respectively. Calculate the ionization energies of an H atom and account for the differences in these three values. 


1
Expert's answer
2021-10-07T03:43:53-0400

Given first IE of Cs=3.89eV (IE=Ionization energy)

First IE of Au=9.23eV

Electronic configuration of Cs(55) : [Xe] 6s1

Electronic configuration of Au(79) : [Xe] 4f14 5d10 6s1

in both elements outermost electron in 6th orbit , so principal quantum number n=6.

According to Bhor theory , Ionization energy of H if electron in nth orbit

"E_n = \frac{13.6}{n^2} \\

n=6 \\

E_n = \frac{13.6}{36} = 0.38 \;eV"

Difference in ionization energy of Cesium(Cs) to Hydrogen(H) =3.89-0.38 =3.51eV

Difference in ionization energy of Gold(Au) to Hydrogen(H)= 9.23-0.38 = 8.85eV

Difference in ionization energy of Gold(Au) to Cesium(Cs)= 9.23-3.89 = 5.34eV